MnO-4+I——-I2+Mn2+
Balance the following reactions using the 1/2 cell method under acidic conditions.
A. 2[5c-+8H+ +Mno-4 ——M2+2+4420
B. 5[2I-1 ——-O-2+2c-
1
Expert's answer
2019-04-09T07:56:49-0400

(A) Mn+ 8 +5 e=M +4O

(B) 2 -2 e=

The A multiply on 2, the В multiply on 5

The balanced equation under acidic condition will be

2 Mn +10+ 16 = 2 M +5 + 8O